### v1

# BA-ORABE: Blockchain-Based Auditable Registered Attribute-Based Encryption With Reliable Outsourced Decryption

Abstract
Attribute-based encryption (ABE) is a generalization of public-key encryption that enables fine-grained access control in cloud services. Recently, Hohenberger et al. (Eurocrypt 2023) introduced the notion of registered ABE, which is an ABE scheme without a trusted central authority. Instead, users generate their own public/secret keys and then register their keys and attributes with a key curator. The key curator is a transparent and untrusted entity and its behavior needs to be audited for malicious registration. In addition, pairing-based registered ABE still suffers the heavy decryption overhead like ABE. A general approach to address this issue is to outsource decryption to a decryption cloud service (DCS). In this work, we propose BA-ORABE, the first fully auditable registered ABE with reliable outsource decryption scheme based on blockchain. First, we utilize a verifiable tag mechanism to achieve verifiability of ciphertext transformation, and the exemptibility which enables the honest DCS to escape from wrong claims is guaranteed by zero knowledge fraud proof under optimistic assumption. Additionally, our system achieves fairness and decentralized outsourcing to protect the interests of all parties and the registration and outsourcing process are transparent and fully auditable through blockchain. Finally, we give formal security analysis and implement and evaluate our scheme on Ethereum to demonstrate its feasibility and efficiency.
keywords:
Registered Attribute-Based Encryption; Outsourced Decryption; Blockchain; NIZK
1 Introduction
The swift progression of cloud computing technologies has led to a significant transformation in data sharing landscape. Nowadays, both individuals and enterprises are increasingly reliant on cloud storage platforms for uploading and disseminating data. However, given that cloud service providers (CSPs) may not always be deemed fully reliable and considering the high sensitivity of much of the data stored in the cloud, ensuring access control within cloud storage is paramount.
In public cloud storage, ciphertext-policy attribute-based encryption (CP-ABE) provides an effective approach to achieve access control[1]. CP-ABE enables fine-grained access control by embedding access structure directly into the encrypted data and associating attributes with decryption keys. This allows data owners to specify access policy in encryption, and decryption is only possible when data user’s attributes match the access policy defined in the ciphertext.
However, Attribute-Based Encryption(ABE) upends the traditional trust model of public-key encryption by relying on a single trusted authority that maintains a long-term master secret key to issue decryption keys[2]. If the central authority is compromised and its secret key is obtained by an adversary, the adversary would be able to decrypt all ciphertexts within the system. The proposal of decentralized ABE still relies on multiple authorities, and therefore, it fails to completely address this issue[3]. To address the key-escrow problem, Hohenberger et al.[4] proposed a registered ABE scheme without any authority that allows users to generate secret keys on their own and then register the associated public key with a key curator along with their attributes. The key curator combines the public keys of different users into a single master public key. Users may need to obtain auxiliary decryption keys from the key curator, which they then combine with their own secret keys to decrypt. The key curator is transparent and does not maintain secrets, allowing for public auditing, as both registration and update algorithms are deterministic. Although a malicious key curator could register attributes that allow the decryption of all ciphertexts, such behavior can always be detected by an external auditor. By utilizing a decentralized, immutable ledger, blockchain[5] can provide auditable records of all operations performed by the key curator. This ensures that any malicious behavior, such as the registration of unauthorized attributes or wrong computation, can be detected by public verifier. The tamper-resistant nature of blockchain enhances the integrity of the system by enabling the verification and traceability of all interactions without reliance on a central authority.
Alternatively, both pairing-based ABE[2, 6, 7] and registered ABE[4, 8] suffer from high decryption overhead, which poses a significant challenge for lightweight devices. To address this issue in ABE, Green et al. [9] proposed an outsourced decryption scheme for attribute-based encryption(OABE), which significantly reduces the computational overhead for users without revealing sensitive information about the original data. In their scheme, user generates a transformation key (TK) and a retrieve key (RK) from the attribute private key (SK) and sends the TK and ciphertext (CT) to the decryption cloud server (DCS) for partial decryption, which transforms the ciphertext into an ElGamal ciphertext [10]. The user then uses the RK to decrypt the ElGamal ciphertext and retrieve the plaintext.
Since the DCS may not perform the computation honestly to save computation resources, many verifiable OABE schemes have been proposed in the literature [11, 12, 13, 14, 15]. However, many previous schemes fail to achieve the properties of exemptibility and fairness. The exemptibility property ensures that the DCS is not falsely accused of malicious behavior if it returns a correct transformed ciphertext, while the fairness property ensures that the DCS gets paid if and only if it returns a correct result. Cui et al. [16] propose a functional encryption with payable outsourced decryption (FEPOD) scheme based on blockchain[5], which achieves these two properties without relying on a trusted authority. Ge et al. [17] eliminates redundant information while achieving exemptibility and fairness, but outsources heavy computation to smart contracts, which will cause high gas usage and performance bottlenecks.
Despite the success of OABE, to the best of our knowledge, there is no outsourcing decryption scheme has been proposed for registered ABE. This gap motivates our work, as we aim to develop first outsourced decryption scheme for registered attribute-based encryption(ORABE) which not only maintains the security and privacy of the data but also achieves verifiability, exemptibility, fairness and decentralized outsourcing.
1.1 Motivation and Contribution
While registered ABE tackles the key-escrow problem upon traditional ABE by removing central authority, it requires robust auditing to ensure the integrity of key registration and update processes, as the transparent key curator still plays a critical role in the system. Additionally, it still suffers from high decryption overhead like ABE, making it unsuitable for lightweight devices. Many verifiable OABE schemes have been proposed to address this issue on ABE and several OABE schemes[16, 17] achieve exemptibility but outsource heavy computation to blockchain, leading to high gas fees and performance issues. To the best of our knowledge, there is no outsource decryption scheme currently existing for registered ABE. This gap motivates our work to develop the first outsourced decryption scheme for registered ABE, which aims to reduce decryption overhead while ensuring security, verifiability, exemptibility, and fairness based on blockchain. In summary, the contributions of this paper are threefold.
  • We propose the first fully auditable registered attribute-based encryption with outsourced decryption scheme based on blockchain, which reduces the decryption overhead to only one group exponentiation operation without pairing operation.
  • We achieve verifiability of transform ciphertext by utilizing a verifiable tag mechanism[13] and achieve exemptibility which enables the honest DCS to escape from wrong claims through NIZK fraud proof under optimistic assumption. Besides, our scheme also achieves fairness and decentralized outsourcing, and the payment can be processed by cryptocurrency based on blockchain.
  • We implement a concrete auditable registered ABE scheme with reliable and payable outsourced decryption and evaluate feasibility and performance on Ethereum.
1.2 Related Work
1.2.1 Outsource Decryption Attribute Based Encryption
Attribute-Based Encryption (ABE) was first introduced as fuzzy identity-based encryption by Sahai and Waters in [18]. ABE schemes are generally classified into two categories: Ciphertext-Policy ABE (CP-ABE) and Key-Policy ABE (KP-ABE) [6], depending on whether the access policy is embedded in the ciphertext or the user’s private key. A user can decrypt only if the set of attributes satisfies the access policy of ciphertext. However, the decryption in the ABE is too expensive, and this problem is especially acute for resource limited devices such as mobile devices, which greatly hinders its real-world adoption. To mitigate this issue, Green et al. [9] proposed outsourcing the majority of the decryption to third party without revealing the plaintext. The pioneer work[9] can append a hash of ciphertext randomness as tag to check the correctness of transformation. Consequently, Lai et al. [11] extend the original ABE model with outsourced decryption to achieve verifiability. Their approach involved attaching an additional ciphertext of a random message and a tag to each original ciphertext which double the size of the original ABE ciphertext and the transformation cost. Based on the scheme [11], Qin et al. [13] and Lin et al. [14] aimed to enhance the efficiency of OABE. Qin et al.[13] managed to reduce the ciphertext size on both the server and client sides by half and achieved faster decryption than the scheme proposed in [11]. They also introduced a method to transform any ABE scheme with outsourced decryption into one that supports verifiable outsourced decryption. Lin et al. [14] proposed a more efficient and generic construction of ABE with verifiable outsourced decryption based on an attribute-based key encapsulation mechanism, a symmetric-key encryption scheme, and a commitment scheme. Furthermore, Li et al. [12] introduced the notion of ABE scheme that support both outsourced encryption and decryption, but only supports the threshold access policy. Following their work, Ma et al. [19] extended it to accommodate any monotonic access policy and established a more stronger security framework compared to [11]. None of the aforementioned schemes can achieve both exemptibility and fairness. To tackle this issue, Cui et al. [16] introduced an payable OABE scheme that leverages blockchain technology to achieve exemptibility and fairness. However, it still introduces redundant information. Ge et al. [17] eliminated redundant information and mitigated heavy computation problem of smart contracts through decomposing to pieces of meta computations, but their solution still results in high gas consumption and performance limitations. Furthermore, due to the lack of support for pairing computation in native Ethereum[20], this scheme is incompatible with Ethereum platform.
1.2.2 Registered Attribute Based Encryption
Registered Attribute-Based Encryption has recently been introduced as a significant extension of ABE, aimed at addressing the key escrow problem [4]. It replaces the trusted key-issuer who maintains a master secret key to issue decryption keys in traditional ABE with a transparent and untrusted key curator to aggregate users’ public keys. In a registered ABE system, users generate their own public/secret key pairs independently, similar to standard public-key encryption. They subsequently register their public keys with a key curator, along with the attributes set. The key curator aggregates these pairs into a compact master public key. A user encrypts a message under an access policy using the master public key to generate a ciphertext. Decryption is only possible for users who possess an attributes set that satisfies the access policy, using a helper decryption key from the key curator and their own secret key. Hohenberger et al. [4] introduced the concept of registered ABE and provided a construction based on assumptions over composite-order pairing groups. Their scheme supports all monotone Boolean formula policies and accommodates a bounded number of users L. However, it relies on a structured common reference string (CRS) whose size scales quadratically with the number of users L and linearly with the size of the attribute universe. Subsequently, several constructions of registered ABE have been proposed, which enhance both the functionality and security of the original scheme[8, 21, 22, 23, 24, 25, 26]. There are some constructions require non-black-box use of cryptography (i.e.,obfuscation-based or witness-encryption-based approaches)[8, 21, 22, 23, 25], and pairing-based constructions that only require black-box use of cryptography[8, 25, 24, 26]. The pairing-based constructions are more efficient, but many of them [4, 25, 24, 26] had a CRS whose size was quadratic in the number of users, and for the schemes that supported general Boolean formulas, also linear in the number of attributes. Grag et al.[8] reducing the CRS size to nearly linear in the number of users L(i.e. $L^{1+o(1)}$) with progression-free sets, which makes the scheme more practical. But the pairing-based construction of [8] still suffers from high decryption overhead. Therefore, we build upon their construction to develop an outsourced decryption scheme aimed at minimizing the decryption burden for users while achieving verifiability, exemptibility, and fairness.
1.3 Organization
The rest of this paper is organized as follows. In Section 2, we briefly revisit the definitions that are relevant to this paper. In Section 3, we describe the system overview of our scheme. In Section 4, we present generic construction of slotted ORABE and ORABE scheme using blockchain and zero-knowledge fraud proof. In section 5, we give formal security analysis of our scheme. In Section 6, we implement the proposed concrete scheme on Ethereum to evaluate its performance. Finally, this paper is concluded in Section 7.
2 Preliminaries
In this section, we review some definitions associated with our scheme.
2.1 Bilinear Map
Let $\mathbb{G},\mathbb{G}_{T}$ are two cyclic groups of prime order $p$, $e$ is the bilinear map $\mathbb{G}\times\mathbb{G}\rightarrow\mathbb{G}_{T}$ and $g$ is the generator of $\mathbb{G}$. The bilinear map has three properties:
  • Bilinearity: $e(g^{a},g^{b})=e(g,g)^{ab}$ holds for all $a,b\in\mathbb{Z}_{p}^{*}$, where $\mathbb{Z}_{p}$ is a field of prime order $p$.
  • Non-degeneracy: $e(g,g)\neq 1$.
  • Computability: $e(g,g)$ can be computed in polynomial time.
2.2 Outsourced Decryption Ciphertext-Policy Attribute-Based Encryption
Definition 2.1.
[…]
  • $\textbf{Setup}(\lambda,U)\rightarrow PK,MSK$. On input a security parameter and a universe description, the setup algorithm outputs the public key PK and a master key MSK.
  • $\textbf{KeyGen}(MSK,S)\rightarrow SK$. On input the master key MSK and a set of attributes S, the key-generation algorithm outputs a secret key SK.
  • $\textbf{Encrypt}(m,PK,(M,\rho))\rightarrow CT$. On input a message m, the public key PK and an LSSS access structure $(M,\rho)$, the encryption algorithm outputs a ciphertext CT.
  • $\textbf{TKGen}(SK)\rightarrow(TK,z)$. On input the secret key SK, the transform key-generation algorithm choose a random value $z\in\mathbb{Z}_{p}^{*}$ and outputs a transform key TK and a retrieve key z.
  • $\textbf{Transform}(CT,TK)\rightarrow CT^{\prime}$. On input the transform key TK and the ciphertext CT, If S does not satisfy the access structure, the transform algorithm outputs $\perp$, else it outputs transform ciphertext $CT^{\prime}$.
  • $\textbf{Decrypt}_{out}(CT^{\prime},z)\rightarrow M$. On input a transform ciphertext $CT^{\prime}$ and a retrieve key z, the decryption algorithm outputs a message M.
2.3 Slotted Registered Attribute-Based Encryption and Registered Attribute-Based Encryption
Definition 2.2.
(Slotted Registered Attribute-Based Encryption[4]). Let $\lambda$ be a security parameter. let $\mathcal{U}$=$\{\mathcal{U}_{\lambda}\}_{\lambda\in\mathbb{N}}$ be a universe of attributes and $\mathcal{P}$=$\{\mathcal{P}_{\lambda}\}_{\lambda\in\mathbb{N}}$ be a set of policies on $\mathcal{U}$. A slotted registered ABE scheme with attribute universe $\mathcal{U}$ and policy space $\mathcal{P}$ is a tuple of efficient algorithms $\prod_{sRABE}=$ $(\textbf{Setup},\textbf{KeyGen},\textbf{IsValid},\textbf{Aggregate},\textbf{Encrypt},\textbf{Decrypt}):$
  • $\textbf{Setup}(1^{\lambda},1^{|\mathcal{U}_{\lambda}|},1^{L})\rightarrow crs$. On input the security parameter $\lambda$, the size of the attribute universe $\mathcal{U}_{\lambda}$ , and the number of slots L, the setup algorithm outputs a common reference string crs.
  • $\textbf{KeyGen}(crs,i)\rightarrow pk_{i},sk_{i}$. On input the common reference string crs, a slot index $i\in[L]$, the key-generation algorithm outputs a public key $pk_{i}$ and a secret key $sk_{i}$ for slot i.
  • $\textbf{IsValid}(crs,i,pk_{i})\rightarrow{0,1}$. On input the common reference string crs, a slot index $i\in[L]$ and a public key $pk_{i}$, the key-validation algorithm outputs a bit $b\in\{0,1\}$ indicating whether $pk_{i}$ is valid or not.
  • $\textbf{Aggregate}(crs,\{(pk_{i},S_{i})\}_{i\in[L]})\rightarrow(mpk,$$\{hsk_{i}\}_{i\in[L]})$. On input the common reference string crs and a list of public keys and the associated attributes $(pk_{1},S_{1})...(pk_{L},S_{L})$, the aggregate algorithm outputs the master public key mpk and a collection of helper decryption keys $hsk_{1},...,hsk_{L}$.
  • $\textbf{Encrypt}(mpk,P,m)\rightarrow ct$. On input the master public key mpk, an access policy $P\in\mathcal{P}_{\lambda}$, and a message $m$, the encryption algorithm outputs a ciphertext ct.
[…]
Definition 2.3.
(Registered Attribute-Based Encryption[4]). Let $\lambda$ be a security parameter. let $\mathcal{U}$=$\{\mathcal{U}_{\lambda}\}_{\lambda\in\mathbb{N}}$ be a universe of attributes and $\mathcal{P}$=$\{\mathcal{P}_{\lambda}\}_{\lambda\in\mathbb{N}}$ be a set of policies on $\mathcal{U}$. A registered ABE scheme with attribute universe $\mathcal{U}$ and policy space $\mathcal{P}$ is a tuple of efficient algorithms $\prod_{RABE}=$ $(\textbf{Setup},\textbf{KeyGen},\textbf{RegPK},\textbf{Encrypt},\textbf{Update},\textbf{Decrypt}):$
  • $\textbf{Setup}(1^{\lambda},1^{|\mathcal{U}_{\lambda}|})\rightarrow crs$. On input the security parameter $\lambda$ and the size of the attribute universe $\mathcal{U}_{\lambda}$, the setup algorithm outputs a common reference string crs.
  • $\textbf{KeyGen}(crs,aux)\rightarrow pk,sk$. On input the common reference string crs and a (possibly empty) state aux, the key-generation algorithm outputs a public key $pk$ and a secret key $sk$.
  • $\textbf{RegPK}(crs,aux,pk,S_{pk})\rightarrow mpk,aux^{\prime}$. On input the common reference string crs, a (possibly empty) state aux, a public key pk, and a set of attributes $S_{pk}\subseteq\mathcal{U}_{\lambda}$, the registration algorithm deterministically outputs the master public key mpk and an updated state aux’.
  • $\textbf{Encrypt}(mpk,P,m)\rightarrow ct$. On input the master public key mpk, an access policy $P\in\mathcal{P}_{\lambda}$, and a message $m$, the encryption algorithm outputs a ciphertext ct.
  • $\textbf{Update}(crs,aux,pk)\rightarrow hsk$. On input the common reference string crs, a state aux, and a public key pk, the update algorithm deterministically outputs a helper decryption key hsk.
  • $\textbf{Decrypt}(sk,hsk,ct)\rightarrow m\cup\{\perp,\textbf{GetUpdate}\}$. On input a decryption key sk, the helper decryption key hsk, and a ciphertext ct, the decryption algorithm either outputs a message m, a special symbol $\perp$ to indicate a decryption failure, or a special flag GetUpdate that indicates an updated helper decryption key is needed to decrypt.
2.4 Non-Interactive Zero Knowledge(NIZK) Protocol for Discrete Logarithm Equality(DLEQ)
Suppose $g,G\in\mathbb{G}_{1}$, $t,T\in\mathbb{G}_{2}$, $|\mathbb{G}_{1}|=|\mathbb{G}_{2}|=p$, and $s\in\mathbb{Z}_{p}.$ $\mathcal{R}_{DLEQ}=\left\{((g,G,t,T),s)|s=log_{g}G=log_{t}T\right\}$, i.e. $((g,G,t,T),s)\in R_{DLEQ}$ if and only if $log_{g}G=log_{t}T$. A sigma protocol for discrete logarithm equality $\prod_{DLEQ}$ is a tuple of probabilistic polynomial-time(PPT) algorithms (Commit, Challenge, Response, Verify) defined as follows.
[…]
  • $\textbf{Response}(\mathbbm{x},\mathbbm{w},r,c)\rightarrow z$. On input public input $\mathbbm{x}=(g,G,t,T)$ and witness $\mathbbm{w}=s$, prover compute $z=r+cs$. Finally, it outputs response $z$ to verifier.
  • $\textbf{Verify}(\mathbbm{x},(C_{1},C_{2},c,z))\rightarrow accept/reject$. On input public input $\mathbbm{x}=(g,G,t,T)$ and $(C_{1},C_{2},c,z)$, verifier checks equations:
    $g^{z}=C_{1}G^{c},t^{z}=C_{2}T^{c}.$
    If both the equations hold, it outputs accept to indicate the discrete logarithm equality of tuple$(g,G,t,T)$ and reject otherwise.
The property of $\prod_{DLEQ}$ sigma protocol:
  • Completeness. For any $(\mathbbm{x},\mathbbm{w})\in\mathcal{R}$, the prover P convinces the verifier V to accept with probability $1$.
  • Special Soundness. From any $\mathbbm{x}$ and any pair of accepting conversations on input $\mathbbm{x}$, $(C_{1},C_{2},c,z),(C_{1},C_{2},c^{\prime},z^{\prime})$ where $c\neq c^{\prime}$, one can efficiently compute $\mathbbm{w}$ such that $(\mathbbm{x},\mathbbm{w})\in\mathcal{R}$.
  • Special Honest-Verifier Zero Knowledge. There exists a polynomial-time simulator M, which on input $\mathbbm{x}$ and a random $c$ outputs an accepting conversation of the form $(C_{1},C_{2},c,z)$, with the same probability distribution as conversations between the honest P, V on input $\mathbbm{x}$.
The above protocol can be made non-interactive in the random oracle model using the Fiat-Shamir transformation[27] to generate challenge. $\prod_{DLEQ-NIZK}$ is a tuple of PPT algorithms (Prove, Verify) defined as follows and $H$ is a hash function.
[…]
  • $\textbf{Verify}(\mathbbm{x},\pi)\rightarrow accept/reject$. On input public input $\mathbbm{x}=(g,G,t,T)$ and proof $\pi=(C_{1},C_{2},c,z)$, the verify algorithm checks equations:
    $c=H(C_{1},C_{2}),g^{z}=C_{1}G^{c},t^{z}=C_{2}T^{c}.$
    If all the equations hold, it outputs accept to indicate the discrete logarithm equality of tuple$(g,G,t,T)$ and reject otherwise.
3 System Overview
In this section, we describe our system architecture, threat model, and security definition.
3.1 System Architecture
Our system consists of seven entities, namely Trusted Setup Parties(TSPs), Key Curator(KC), InterPlanetary File System(IPFS), Data Owners(DOs), Data Users(DUs), Decryption Cloud Server(DCS) Network and Public Verifier(PV). These seven entities are defined as follows:
  • TSPs are responsible for one-time trusted setup through decentralized protocol[28] to generate common reference string crs.
  • KC is responsible for aggregating the public keys registered by users and generating master public key and helper decryption keys for all users.
  • IPFS is a distributed file system to store data without a central server. It stores auxiliary data and ciphertexts in our system.
  • DOs use specified access structure to encrypt data, upload the ciphertext to IPFS, and publish the corresponding ciphertext tag on blockchain.
  • DUs publish outsourced decryption task on blockchain, get transform ciphertext from it, and finish final decryption with secret key. Besides, DUs can generate fraud proof to accuse dishonest DCS.
  • DCS network consists of many servers that compete for tasks. DCSs get ciphertexts from IPFS, request helper decryption key from KC, compute and submit transform ciphertext to blockchain.
  • PV verifies fraud proofs submitted by DUs and publishes the verification result on blockchain.
Figure 1: Overview of happy case.
Fig 1 illustrates the overview of happy case of our scheme, which consists of the following steps:
  • TSPs execute trusted setup ceremony and generate common reference string CRS.
  • DU generates private key and public key locally, then registers their public key($\mathbf{pk}_{u}$) and attributes set to KC.
  • KC firstly verifies DU’s public key($\mathbf{pk}_{u}$). If $\mathbf{pk}_{u}$ is valid, KC computes and updates auxiliary data on IPFS, and publishes new state hash $\mathbf{hash(aux)}$ and public key hash $\mathbf{hash(pk_{u})}$ on blockchain.
  • DO encrypts sensitive data using RABE under Key Encapsulation Mechanism(KEM) and sends ciphertext to IPFS.
  • DO publishes the associated ciphertext tag to blockchain so that DU can check the correctness of transform ciphertext with the tag.
  • DU publishes outsourced decryption task on blockchain.
  • DCS observes outsourced decryption task and downloads corresponding ciphertext and auxiliary data from IPFS.
  • DCS transforms the original ciphertext and submits the transform ciphertext to blockchain.
  • DU gets the result from blockchain and finishes the decryption locally.
  • DCS claims task reward from blockchain after challenge window.
Figure 2: Overview of dispute phase.
After the final decryption, DU could check whether the decrypted plaintext matches the ciphertext tag. If not, DU can enter the dispute phase which involves four more steps as shown in Fig 2:
  • DU generates fraud proof to accuse DCS’s malicious behavior and publish it on blockchain.
  • PV gets fraud proof from blockchain and verifies its validity.
  • PV publishes the verification result on blockchain.
  • Smart contract will return DU’s deposit if the fraud proof is valid.
Note that the dispute phase and public verifier are needed only when one player crashes or attempts to cheat under our optimistic assumption.
3.2 Threat Model
In our scheme, DOs and IPFS are fully trusted. TSPs execute one-time trusted setup through decentralized protocol, the setup is secure as long as any single party participates honestly[28]. KC and PV are semi-honest and transparent, which means they maintain no secret and perform computation honestly, but are curious about secrets. DCSs are malicious, indicating that they may try to obtain task reward by submitting incorrect results deliberately. Considering that DUs intend to propose invalid fraud proof in order not to pay, we assume DU to be malicious as well.
3.3 Security Definition
  • Static Security. Let $\prod_{sORABE}=(\textbf{Setup},\textbf{KeyGen},\textbf{IsValid},\textbf{Aggregate},\textbf{Encrypt},$ Transform, $\textbf{Decrypt}_{out})$ be a slotted outsource decryption registered ABE scheme with attribute universe $\mathcal{U}$ and policy space $\mathcal{P}$. For a security parameter $\lambda$, an adversary $\mathcal{A}$, and a bit $b\in\{0,1\}$, we define the following game $Game_{sec}$ between a challenger C and an adversary $\mathcal{A}$:
    • Setup: C runs Setup algorithm to generate crs, initializes the auxiliary input aux = $\perp$, the initial master public key mpk = $\perp$, a counter ctr = 0 for the number of honest-key-registration queries the adversary has made, and an empty dictionary mapping public keys to registered attribute sets Dict. The challenger gives the crs to $\mathcal{A}$.
    • Query: The challenger responds to the adversary’s key-generation queries on a non-corrupted slot index $i\in N$ with a set of attributes S. The challenger starts by incrementing the counter $ctr=ctr+1$ and samples $pk_{ctr},sk_{ctr}\leftarrow KeyGen(1^{\lambda})$. The challenger replies to $\mathcal{A}$ with $(ctr,pk_{ctr})$. Note that in the static security game, the adversary is not allowed to make any corruption queries.
    • Challenge: The adversary $\mathcal{A}$ chooses two messages $\mu_{0}^{*},\mu_{1}^{*}$ in message space and an access policy $P^{*}$ in policy space. The challenger replies with the challenge ciphertext $ct^{*}\leftarrow\textbf{Encrypt}(mpk,P^{*},\mu_{b}^{*}).$
    • Output: At the end of the game, $\mathcal{A}$ outpus a bit $b^{\prime}\in\{0,1\}$.
    The advantage of the adversary in $Game_{sec}$ is defined as $\textbf{Pr}[b^{\prime}=b]-\frac{1}{2}$.
    Definition 3.1.
    A Registered ABE scheme with outsourced decryption is secure if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
    $\textbf{Pr}[b^{\prime}=b]-\frac{1}{2}\leq negl(\lambda)$
  • Verifiability of outsourced decryption. Verifiability guarantees that the correctness of transform ciphertext can be efficiently checked. It is described by a game $Game_{ver}$ between a challenger C and an adversary $\mathcal{A}$. $Game_{ver}$ consists of the following steps:
    • Setup: C runs Setup gives crs to $\mathcal{A}$.
    • Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to C, C runs RegPK and gives mpk to $\mathcal{A}$.
    • Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},sk^{*},CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}\}$, where S* satisfies $(M^{*},\rho^{*})$, $CT^{*}$ is an original ciphertext, $sk^{*}$ is the secret key of $\mathcal{A}$ and $CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}$ are two transformed ciphertexts of $CT^{*}$. $\mathcal{A}$ wins the game if:
      $Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq\perp\cap$
      $Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})\neq\perp\cap$
      $Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})$
    The advantage of the adversary in $Game_{ver}$ is defined as $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{ver}]$.
    Definition 3.2.
    A Registered ABE scheme with outsourced decryption is verifiable if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
    $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{ver}]\leq negl(\lambda)$
  • Exemptibility of outsourced decryption. Exemptibility guarantees that the user can not falsely accuse DCS of returning wrong transform ciphertext if it is not in the case. It is described by a game $Game_{exe}$ between a challenger C and an adversary $\mathcal{A}$. $Game_{exe}$ consists of the following steps:
    • Setup: C runs Setup and gives crs to $\mathcal{A}$.
    • Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to C, C runs RegPK and gives mpk and aux to $\mathcal{A}$.
    • Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},CT^{*^{\prime}},hsk*,(sk_{1}^{*},\pi_{1}),(sk_{2}^{*},\pi_{2}),T_{i}^{*}\}$, where S* satisfies $(M^{*},\rho^{*})$, $CT^{*}$ is an original ciphertext, $CT^{*^{\prime}}$ is the transform ciphertext of $CT^{*}$, $hsk*,sk_{1}^{*},sk_{2}^{*}$ are the helper decryption key and secret keys, $T_{i}^{*}=g^{sk^{*}}$ is part of public key of $\mathcal{A}$ and $\pi_{1},\pi_{2}$ are the corresponding NIZK proof for secret keys. $\mathcal{A}$ wins the game if:
      $hsk^{*}\leftarrow Update(crs,aux,pk^{*})\cap$
      $CT^{*^{\prime}}\leftarrow Transform(hsk^{*},CT^{*})\cap$
      $1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{1}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*}))\cap$
      $1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{2}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*}))\cap$
      $Decrypt_{out}(sk_{1}^{*},CT^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk_{2}^{*},CT^{*^{\prime}},CT^{*})$
    The advantage of the adversary in $Game_{exe}$ is defined as $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{exe}]$.
    Definition 3.3.
    A Registered ABE scheme with outsourced decryption is exempt if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
    $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{exe}]\leq negl(\lambda)$
4 Construction
We first present slotted registered ABE and registered ABE scheme with outsourced decryption based on registered ABE scheme in [8]. Then we describe a NIZK fraud proof mechanism under optimistic assumption to achieve exemptibility. Finally, we present the concrete smart contract design for our payable outsourced decryption scheme.
4.1 Verifiable Outsourced Decryption Slotted Registered ABE
Our proposed slotted registered ABE with outsourced decryption(sORABE) scheme is as follows.
  • $\textbf{Setup}(1^{\lambda},1^{L})\rightarrow crs$. On input the security parameter $\lambda$ and the number of slots L, the setup algorithm sampling $\mathcal{G}=(\mathbb{G},\mathbb{G}_{T},g,e)\leftarrow PrimeOrderGen(1^{\lambda})$. And then it constructs:
    • Let $\mathcal{D}=\{d_{i}\}_{i\in[L]}$ be an efficiently-computable progression-free and double-free set of size L. We define $f(i,j):=d_{i}+d_{j}$ and the set $\mathcal{E}$ of all distinct pairwise sums of elements in $\mathcal{D}$:
      =$\mathcal{E}=\{f(i,j)|i,j\in[L]:i\neq j\}$
      let $d_{max}=3\cdot\text{max}(\mathcal{D})$.
[…]
      $A_{i}=g^{t_{i}},B_{i}=g^{\alpha}h^{t_{i}},P_{i}=g^{\delta_{i}},U_{i}=g^{bt_{i}}.$
      For each $z\in\mathcal{E}$, let $W_{z}=g^{ba^{z}}$.
    • Finally let $Z=e(g,g)^{\alpha}$ and it outputs the common reference string $crs$
      $crs=(\mathcal{G},Z,g,h,\{(A_{i},B_{i},P_{i},U_{i})\}_{i\in[L]},\{W_{z}\}_{z\in\mathcal{E}})$
[…]
    $T_{i}=g^{r_{i}},Q_{i}=P_{i}^{r_{i}}.$
    Then for each $j\neq i$, it computes the cross terms $V_{j,i}=A_{j}^{r_{i}}$. Finally, it outputs the public key $pk_{i}$ and the secret key $sk_{i}$
    $pk_{i}=(T_{i},Q_{i},\{V_{j,i}\}_{j\neq i})\quad\text{and}\quad sk_{i}=r_{i}.$
  • $\textbf{IsValid}(crs,i,pk_{i})\rightarrow\{0,1\}$. On input the common reference string crs, a slot index $i\in[L]$ and a public key $pk_{i}$, the key-validation algorithm first checks that each of the components in $pk_{i}$ is a valid group element, and then checks
    $e(T_{i},P_{i})=e(g,Q_{i})$
    Next, for each $j\neq i$, the algorithm checks that
    $e(g,V_{j,i})=e(T_{i},A_{j})$
    If all checks pass, it outputs 1; otherwise, it outputs 0.
  • $\textbf{Aggregate}(crs,(pk_{1},S_{1}).(pk_{L},S_{L}))\rightarrow(mpk,hsk_{1},...,hsk_{L})$. On input the common reference string crs and a list of public keys and the associated attributes $(pk_{1},S_{1})...(pk_{L},S_{L})$, the aggregate algorithm starts by computing the attribute-independent public key $\hat{T}$ and the attribute-independent slot key $\hat{V_{i}}$ for each $i\in[L]$:
    $\hat{T}=\prod_{j\in[L]}T_{j}\quad\text{and}\quad\hat{V}_{i}=\prod_{j\neq i}V_{i,j}$
    Next, for each attribute $w\in\mathcal{U}_{\lambda}$, it computes the attribute-specific public key $\hat{U_{w}}$ and the attribute-specific slot key $\hat{W_{i,w}}$ for each $i\in[L]$:
    $\hat{U}_{w}=\prod_{j\in[L]:w\not\in S_{j}}U_{j}\quad\text{and}\quad\hat{W}_{i,w}=\prod_{j\neq i:w\not\in S_{j}}W_{f(i,j)}.$
    Finally, it selects three hash functions $H_{0},H_{1},H_{2}$ and outputs the master public key mpk and the slot-specific helper decryption keys $hsk_{i}$ where
    $mpk=(\mathcal{G},Z,g,h,\hat{T},\hat{U}_{w},H_{0},H_{1},H_{2})$
    $hsk_{i}=(mpk,i,S_{i},A_{i},B_{i},\hat{V}_{i},\{\hat{W}_{i,w}\}_{w\in\mathcal{U}_{\lambda}})$
  • $\textbf{Encrypt}(mpk,(\textbf{M},\rho),m)\rightarrow ct$: On input the master public key mpk, the policy $(\textbf{M},\rho)$ and a message m, it first chooses a session key $\mu\in\mathbb{G}_{T}$ and compute $t_{2}$ as symmetric key. Enc is a symmetric encryption algorithm.
    $t_{1}=H_{0}(\mu),t_{2}=H_{1}(\mu)$
[…]
    $C=\textbf{Enc}_{t_{2}}(m),tag=H_{2}(t_{1}||C)$
    $C_{1}=\mu*Z^{s},C_{2}=g^{s}$
    $C_{3,k}=h_{2}^{m_{k}^{T}v}\hat{U}_{\rho(k)}^{-s_{k}},C_{4,k}=g^{s_{k}}$
    $C_{5}=(h_{1}\hat{T}^{-1})^{s}$
    Finally, it outputs the ciphertext $ct$ where
    $ct=((\textbf{M},\rho),C,C_{1},C_{2},\{C_{3,k},C_{4,k}\}_{k\in[K]},C_{5},tag)$
  • $\textbf{Transform}(hsk,ct)$rightarrow ct’: On input the ciphertext $ct$ and the helper decryption key $hsk$. The transform algorithm proceeds as follows:
    • If the set of attributes $S_{i}$ is not authorized by $(\textbf{M},\rho)$, then the transform algorithm outputs $\perp$.
    • Otherwise, let $I=\{k\in[K]:\rho(k)\in S_{i}\}$ be the indices of the rows of M associated with the attributes $S_{i}\subseteq\mathcal{U}_{\lambda}$.
    • Let $\textbf{M}_{S_{i}}$ be the matrix formed by taking the subset of rows in M indexed by $I$. Since $S_{i}$ is authorized, let $\omega_{S_{i}}\in\mathbb{Z}_{p}^{|I|}$ be a vector such that $\omega_{S_{i}}^{\top}\textbf{M}_{S_{i}}=\textbf{e}_{1}^{\top}$.
    • Then compute:
      $C_{1}^{\prime}=\frac{C_{1}}{e(C_{2},B_{i})}\cdot e(C_{5},A_{i})\cdot e(C_{2},\hat{V_{i}})\cdot\\ \prod_{1\leq j\leq|I|}\big{(}e(C_{3,k_{j}},A_{i})\cdot e(C_{4,k_{j}},\hat{W}_{i,\rho(k_{j})})\big{)}^{\omega_{S_{i,j}}}$
      $C_{2}^{\prime}=e(C_{2},A_{i})$
    • Finally, it outputs the transform ciphertext $ct^{\prime}$ where
      $ct^{\prime}=(C_{1}^{\prime},C_{2}^{\prime})$
  • $\textbf{Decrypt}_{out}(sk,ct^{\prime},ct)$rightarrow $m/\perp$: On input the original ciphertext $ct$, the transform ciphertext $ct^{\prime}$,and secret key sk, the decryption algorithm first compute:
    $\mu^{\prime}=C_{1}^{\prime}\cdot C_{2}^{\prime r_{i}}$
    Next, check the following equation:
    $tag=H_{2}(H_{0}(\mu^{\prime})||C).$
    If it does not hold, the algorithm output $\perp$ which indicates the outsourced decryption result is not correct and DU need to generate a proof to accuse DCS. Otherwise, it uses the symmetric decryption algorithm Dec to recover message
    $m=\textbf{Dec}_{H_{1}(\mu^{\prime})}(C).$
    and outputs message $m$.
4.2 Verifiable Outsourced Decryption Registered ABE
Grag et al.[8] use slotted registered ABE to registered ABE transformation proposed in [4]. We also use the same transformation to construct our outsourced decryption registered ABE scheme from the outsourced decryption slotted registered ABE scheme. Our solution can also be regarded as designing an outsourced decryption scheme based on the registered ABE construction in Appendix C of [8].
  • $\textbf{Setup}(1^{\lambda},1^{L})\rightarrow crs$. On input the security parameter $\lambda$ and the number of slots $L=2^{l}$, the setup algorithm runs the setup algorithm for $l+1$ copies of the slotted outsourced decryption registered ABE(sORABE) scheme. For each $k\in[0,l]$, it samples $crs_{k}\leftarrow\textbf{sORABE.Setup}(1^{\lambda},1^{L})$ and outputs $crs=(crs_{0},...,crs_{l})$.
  • $\textbf{KeyGen}(crs,aux)\rightarrow pk,sk$. On input the common reference string crs, and the auxiliary data $\textbf{aux}=(ctr,Dict_{1},Dict_{2},mpk)$, the key-generation algorithm generates a public/secret key-pair for each of the $l+1$ schemes. For each $k\in[0,l]$, let $i_{k}=(ctr\quad mod\quad 2^{k})+1\in[2^{k}]$ be a slot index for the k-th scheme, and samples $(pk_{k},sk_{k})\leftarrow\textbf{KeyGen}(crs_{k},i_{k}).$ Output $pk=(ctr,pk_{0},...,pk_{l})$ and $sk=(ctr,sk_{0},...,sk_{l}).$
  • $\textbf{RegPK}(crs,aux,pk,S_{pk})\rightarrow mpk^{\prime},aux^{\prime}$. On input the common reference string $crs=(crs_{0},...,crs_{l})$, the auxiliary data $aux=(ctr_{aux},Dict_{1},Dict_{2},mpk)$, where $mpk=(ctr_{aux},mpk_{0},...,mpk_{l})$, public key $pk=(ctr_{pk},pk_{0},...,pk_{l})$, and associated attributes $S_{pk}$, the registration algorithm executes as follows:
    • For each $k\in[0,l]$, let $i_{k}=(ctr_{aux}\quad mod\quad 2^{k})+1\in[2^{k}]$ be the slot index for the k-th scheme.
[…]
    • Finally, the registration algorithm outputs $mpk^{\prime}=(ctr_{aux}+1,mpk^{\prime}_{0},...,mpk^{\prime}_{l})$ and $aux^{\prime}=(ctr_{aux}+1,Dict_{1},Dict_{2},mpk^{\prime})$.
  • $\textbf{Encrypt}(mpk,(\textbf{M},\rho),m)\rightarrow ct$: On input the master public key $mpk=(ctr,mpk_{0},...,mpk_{l})$, the access policy $(\textbf{M},\rho)$ and a message m, for each $k\in[0,l]$, the encryption algorithm computes $ct_{k}\leftarrow\textbf{sORABE.Encrypt}(mpk_{k},(\textbf{M},\rho),m)$, if $mpk=\perp$, then it sets $ct_{k}=\perp$. It outputs $ct=(ctr,ct_{0},...,ct_{l})$.
  • $\textbf{Update}(crs,aux,pk)\rightarrow hsk$: On input the common reference string $crs=(crs_{0},...,crs_{l})$, the auxiliary data $aux=(ctr_{aux},Dict_{1},Dict_{2},mpk)$, and a public key $pk=(ctr_{pk},pk_{0},...,pk_{l})$, if $ctr_{pk}\leq ctr_{aux}$, for each $k\in[0,l]$, it sets $hsk_{k}=Dict_{2}[ctr_{pk}+1,k]$ and output $hsk=(hsk_{0},...,hsk_{l})$. Otherwise, it outputs $\perp$.
  • $\textbf{Transform}(hsk_{k},ct_{k})$rightarrow ct’: On input the ciphertext $ct_{k}$ and the helper decryption key $hsk_{k}$. The algorithm computes $ct^{\prime}\leftarrow\textbf{sORABE.Transform}(hsk_{k},ct_{k})$ and returns ct’.
[…]
4.3 NIZK Fraud Proof Mechanism
Fraud proof, also known as fault proof, is a proof submitted by a challenger that challenges the previous result under optimistic assumption. If the decrypt algorithm outputs $\perp$, it indicates that the transform ciphertext is not correct and DU needs to generate fraud proof to accuse DCS and get the deposit back. The fraud proof generation and verification based on the Discrete Logarithm Equality(DLEQ) NIZK proof is defined as follows:
[…]
    $\pi=(\pi^{\prime},vk).$
  • $\textbf{Verify}(\pi,ct^{\prime},mpk,tag)\rightarrow 0/1$: On input the fraud proof $\pi$, the transform ciphertext $ct^{\prime}$, the master public key mpk, and the ciphertext tag. The verify algorithm first runs $\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi^{\prime},(C^{\prime}_{2},vk,g,T_{i}))$ to check the correctness of $vk$. If the DLEQ NIZK proof is invalid, it outputs 0. Otherwise, it computes $\mu^{\prime\prime}=C_{1}^{\prime}\cdot vk$ and checks the equation $tag=H_{2}(H_{0}(\mu^{\prime\prime})||C).$ If the equation does not hold, it outputs 1 to indicate the fraud proof generated by DU is valid, and 0 otherwise.
Unfortunately, this verify algorithm can not be executed by Ethereum smart contract because it does not support $\mathbb{G}_{t}$ computation of any elliptic curve now. So, we introduce a semi-honest and transparent public verifier to verify the fraud proof and publish verification result on blockchain. Ge et al. [17] added precompile contract to native Ethereum to support pairing and $\mathbb{G}_{t}$ computation on their own blockchain. If reduce the reliance on native Ethereum security, we can integrate their approach to eliminate the role for semi-honest public verifier and enable the verification of fraud proof directly through smart contracts. In this work, we make a trade-off and prioritize the security of native Ethereum. Note that in the registered ABE scheme, this role can also be played by KC, who is semi-honest and transparent.
4.4 Smart Contract
Fig 3 illustrates the BA-ORABE smart contract of our scheme, which manages the whole procedure of user registration, ciphertext tag publication, outsourced decryption, and public verification of fraud proofs. The addresses of KC and PV are recorded in contract once it is deployed.
In our ORABE scheme, each DU first registers itself to KC. On receiving a $\mathbf{pk}_{u}$ from a DU, KC allocates a suitable $ctr$ to DU and updates $aux$ data. To make sure that KC’s behavior is fully auditable, KC should PublishState when it finishes registering DU. The state data indicates the input and output of the registration, which consists of $ctr$, hash($\mathbf{pk}_{u}$), and hash($aux$).
Ciphertexts are stored in IPFS, and DO should PublishTag after uploading ciphertext. Smart contract will record the index of ciphertext in IPFS and its relative tag. The index is used to stipulate which ciphertext to be decrypted, while the tag is used by DU to check the correctness of partially decrypted result and used by PV to verify the validity of fraud proof.
Regarding outsourced decryption, DU first publishes an outsourced decryption task by CreateTask. Note that each task contains some cryptocurrency as task reward. DCS retrieves the corresponding ciphertext from IPFS, performs transformation(partial decryption), and SubmitResult to the blockchain. DU completes the decryption locally upon receiving the result. In the happy case, DCS could ClaimReward after the challenge window. While in the dispute phase, DU could generate a fraud proof and PublishFraudProof during the challenge window. PV verifies the proof, and PublishVerificationResult on blockchain. The auxiliary function _DisposeReward is invoked by PublishFraudProof. It automatically sends the task reward back to the DU if the proof is valid, and to the DCS otherwise.
In pseudocode, msg.sender is the invoker of the contract function, KCstates is the list of all states published by KC, Tags is the map associating the data index in IPFS and the data tag of the same ciphertext, and Tasks is the map associating task identifiers with of all the tasks created. $T_{now}$ denotes the current time; $T_{p}$ denotes challenge window. Transfer($U$, $x$) is a function transferring cryptocurrency of amount $x$ to user $U$.
Figure 3: Contract design for BA-ORABE.
5 Security Analysis
5.1 Static Security
Theorem 5.1.
Suppose the construction 4.3 of [8] is statically secure, then our sRABE with outsource decryption scheme 4.1 is also statically secure under KEM setting.
Proof 5.2.
Suppose there exists a PPT adversary $\mathcal{A}$ that can attack our scheme in the CPA security model with advantage $\epsilon$, then we can build a simulator $\mathcal{B}$ that can break the CPA security of the registered ABE scheme construction 4.3 in [8] with advantage $\epsilon$.
Setup. Challenger $\mathcal{C}$ runs Setup algorithm to generate crs, initializes the auxiliary input aux = $\perp$, the initial master public key mpk = $\perp$, a counter ctr = 0 for the number of honest-key-registration queries the adversary has made, and an empty dictionary mapping public keys to registered attribute sets Dict. The simulator $\mathcal{B}$ got crs from the challenger $\mathcal{C}$ and returns it to the adversary $\mathcal{A}$.
Query. The simulator $\mathcal{B}$ answers the adversary $\mathcal{A}$’s key-generation queries on a non-corrupted slot index $i\in N$ with a set of attributes S. follows:
  • The simulator $\mathcal{B}$ passes it to the challenger $\mathcal{C}$ and $\mathcal{C}$ starts by incrementing the counter $ctr=ctr+1$ and samples $pk_{ctr},sk_{ctr}\leftarrow KeyGen(1^{\lambda})$. The challenger replies to $\mathcal{B}$ with $(ctr,pk_{ctr})$. $\mathcal{B}$ returns the result that answered by $\mathcal{C}$ to $\mathcal{A}$.
Challenge. $\mathcal{A}$ submits a encapsulated key pair $(K_{0}^{*},K_{1}^{*})\in\{0,1\}^{2\times k}$ to $\mathcal{B}$, and $\mathcal{B}$ sends them to the challenger $\mathcal{C}$. The challenger $\mathcal{C}$ generate the challenge ciphertext pair $(CT_{0}^{*},CT_{1}^{*})$ corresponds to $(K_{0}^{*},K_{1}^{*})$ and sends it to $\mathcal{B}$.
Guess. $\mathcal{B}$ passes ciphertext pair $(CT_{0}^{*},CT_{1}^{*})$ to $\mathcal{A}$ and gets the response bit 0 or 1. $\mathcal{B}$ outputs $\mathcal{A}$’s response as its guess.
Note that, the simulation of $\mathcal{B}$ is perfect if the above game does not abort and the construction 4.3 of [8] was proven statically secure under the q-set-consistent bilinear Diffie-Hellman exponent assumption. The advantage of the simulator that wins the CPA game is the same as a adversary that breaks the CPA security of the construction in [8]. Thus, our proposed scheme 4.1 is statically secure.
Theorem 5.3.
Suppose our sRABE with outsource decryption scheme 4.1 is statically secure, then our RABE with outsource decryption scheme 4.2 is also statically secure.
Proof 5.4.
While [4] only analyzing the transformation as applied to fully secure slotted registered ABE scheme, the same analysis also applies to a statically-secure slotted registered ABE scheme[8]. So our transformed registered ABE scheme inherits the same security properties as the slotted scheme.
5.2 Verifiability
Theorem 5.5.
Assume that $H_{0}$, $H_{2}$ are collision-resistant hash functions, the RABE scheme with outsourced decryption is verifiable.
Proof 5.6.
Assume that there exists an adversary $\mathcal{A}$ that can break the verifiability of outsourced decryption, then we construct a simulator $\mathcal{S}$ can break the collision-resistance of hash functions $H_{0},H_{2}$. Given two challenging hash functions $H_{0}^{*},H_{2}^{*}$. $\mathcal{S}$ interacts with $\mathcal{A}$ as follows:
Setup: $\mathcal{S}$ runs Setup gives crs to $\mathcal{A}$.
Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to $\mathcal{S}$, $\mathcal{S}$ runs RegPK and gives mpk to $\mathcal{A}$.
Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},sk^{*},CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}\}$, where S* satisfies $(M^{*},\rho^{*})$. $\mathcal{A}$ wins the game if:
$Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq\perp\cap$
$Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})\neq\perp\cap$
$Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})$
And then $\mathcal{S}$ can get:
$H_{2}^{*}(H_{0}^{*}(\mu_{1}^{*})||C)=tag=H_{2}^{*}(H_{0}^{*}(\mu_{2}^{*})||C),\mu_{1}^{*}\neq\mu_{2}^{*}.$
If $H_{0}^{*}(\mu_{1}^{*})=H_{0}^{*}(\mu_{2}^{*})$, then $\mathcal{S}$ can break the collision-resistance of $H_{0}^{*}$, else $\mathcal{S}$ can break the collision-resistance of $H_{2}^{*}$.
Table 1: Comparison of Pairing-Based RABE Schemes
[…]
[4] $|\mathcal{U}|L^{2}$ $|\mathcal{U}|L^{2}$ $|\mathcal{U}|$ $|\mathcal{U}|$ $|\mathcal{P}|$ $|\mathcal{P}|(\mathbb{P}+\mathbb{E}_{T})$ Composite-order Adaptive
[24] $|\mathcal{U}|L^{2}$ $|\mathcal{U}|L^{2}$ $|\mathcal{U}|$ $|\mathcal{U}|$ $|\mathcal{P}|$ $|\mathcal{P}|(\mathbb{P}+\mathbb{E}_{T})$ Prime-Order Adaptive
[26] $L^{2}$ $L^{2}$ $\delta^{2}$ $\delta^{2}$ $|\mathcal{P}|$ $|\mathcal{P}|(\mathbb{P}+\mathbb{E}_{T})$ Prime-Order Adaptive
[8] $L^{(1+o(1))}$ $|\mathcal{U}|L$ $|\mathcal{U}|$ $|\mathcal{U}|$ $|\mathcal{P}|$ $|\mathcal{P}|(\mathbb{P}+\mathbb{E}_{T})$ Prime-Order Static
Ours $L^{(1+o(1))}$ $|\mathcal{U}|L$ $|\mathcal{U}|$ $|\mathcal{U}|$ $|\mathcal{P}|$ $\mathbb{E}_{T}$ Prime-Order Static
  • Note: crs denotes the size of the common reference string, st denotes the auxiliary information, mpk denotes the size of the master public key, hsk denotes the size of the helper decryption key, and ct denotes the size of the ciphertext. L denotes numbers of users, $\mathcal{U}$ is attribute universe , and $\mathcal{P}$ is policy. In scheme [26], $\delta=max_{i\in[L]}|S_{i}|$ for a set of registered key attribute sets.The decryption cost contains pairing operation $\mathbb{P}$, $\mathbb{G}_{T}$ exponentiation operation $\mathbb{E}_{T}$, and $\mathbb{G}$ exponentiation operation $\mathbb{E}_{G}$. A scheme is statically secure if the adversary is not allowed to make any corruption queries in the security game Definition 3.1 and that it is adaptively secure otherwise.
5.3 Exemptibility
Theorem 5.7.
Suppose that the soundness of $\prod_{DLEQ-NIZK}$ protocol holds, the RABE scheme with outsourced decryption is exempt.
Proof 5.8.
Assume that there exists an adversary $\mathcal{A}$ that can break the exemptibility of outsourced decryption, then we construct a simulator $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}$ protocol. Given challenging protocol $\prod_{DLEQ-NIZK}^{*}$. $\mathcal{S}$ interacts with $\mathcal{A}$ as follows:
Setup: $\mathcal{S}$ runs Setup gives crs to $\mathcal{A}$.
Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to $\mathcal{S}$, $\mathcal{S}$ runs RegPK and gives mpk and aux to $\mathcal{A}$.
Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},CT^{*^{\prime}},hsk*,(sk_{1}^{*},\pi_{1}),(sk_{2}^{*},\pi_{2}),T_{i}^{*}\}$, where S* satisfies $(M^{*},\rho^{*})$. $\mathcal{A}$ wins the game if:
$hsk^{*}\leftarrow Update(crs,aux,pk^{*})\cap$
$CT^{*^{\prime}}\leftarrow Transform(hsk^{*},CT^{*})\cap$
$1\leftarrow\prod_{DLEQ-NIZK}^{*}.\textbf{Verify}(\pi_{1}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*}))\cap$ $1\leftarrow\prod_{DLEQ-NIZK}^{*}.\textbf{Verify}(\pi_{2}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*}))\cap$ $Decrypt_{out}(sk_{1}^{*},CT^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk_{2}^{*},CT^{*^{\prime}},CT^{*}).$ And then $\mathcal{S}$ can get
$sk_{1}^{*}\neq sk_{2}^{*}\ \cap$
$(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*})\in R_{DLEQ}\ \cap$
$(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*})\in R_{DLEQ}$
Thus $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}^{*}$ protocol.
5.4 Fairness
The fairness of our scheme ensures that the DCS gets paid if and only if it returns a correct result. The DU must deposit tokens into the smart contract to submit the outsourced decryption task. If the DCS submits the correct result, it will receive the reward at the end of the challenge window, which is ensured by exemptibility. Otherwise, DU will send valid fraud proof and get the deposit back, and DCS get nothing but pay for the transaction gas fee for malicious behavior. Thus the proposed scheme achieves the fairness property as long as the exemptibility and Ethereum consensus protocol is guaranteed.
Table 2: Comparison With Outsourced Decryption ABE Schemes
Schemes confidentiality verifiability exemptibility fairness decentralized outsourcing
[11]
[12]
[13]
[14]
[19]
[17]
Our Scheme
6 Performance and Evaluation
6.1 Theoretical Analysis
Table 1 shows the comparison of some existing Registered ABE schemes and our scheme. Since our scheme is based on [8], the parameters, except for the decryption cost, are identical to those of [8]. And the parameter sizes are better than other schemes. Besides, our outsourced decryption significantly reduces the final decryption cost to a constant level, requiring only one group exponentiation operation and one group multiplication operation.
Table 2 shows the functional comparison of the previous OABE schemes and our scheme in terms of confidentiality, verifiability, exemptibility, fairness, and decentralized outsourcing. Almost all schemes achieve confidentiality and verifiability. Only our scheme and [17] achieves exemptibility, fairness, and decentralized outsourcing property.
6.2 Experimental Evaluation
We conducted experiments in three aspects: slotted registered ABE computation cost, zero-knowledge fraud proof cost, and smart contract gas usage. In our experiment, the access policy is set as an AND gate of attributes and its size varies from 5 to 100, with a step size of 5.
6.2.1 Outsourced Decryption Slotted Registered ABE(sORABE) vs. Slotted Registered ABE(sRABE)
As we use the slotted registered ABE to registered ABE transformation[4], which constructs registered ABE through multiple slotted registered ABE instances, We only compared the basic slotted scheme here. We conducted experiment and performance analysis on the computation cost comparison between sRABE scheme in [8] and our sRABE with outsourced decryption scheme. We implemented them in Python language within the Charm framework. Our benchmark uses SS512 and SS1024 elliptic curves, SHA-256 hash function, and is executed on an Intel(R) Core(TM) i9-12900H CPU and 4 GB RAM running Ubuntu 20.04 LTS 64-bit and Python 3.8.10. For each algorithm, we repeat 1,000 times to get the average execution time.
As shown in Fig 4(a), the decryption time of sRABE scheme in [8] increases linearly with the numbers of attributes while our scheme with outsource decryption has constant decryption time and is 108$\times$-1585$\times$ in SS512 curve and 175$\times$-2647$\times$ in SS1024 curve lower than [8] in attribute numbers from 5 to 100, which is shown in Fig 4(b). The original and transform ciphertext size of our scheme are shown in Fig 4(c) and Fig 4(d) respectively. Note that we only consider RABE ciphertext and the symmetric encryption ciphertext is not included. The original ciphertext linearly increases with the numbers of attributes like [8] and the transform ciphertext contains only two $\mathbb{G}_{T}$ elements, which is about 0.25kb in SS512 curve and 0.5kb in SS1024 curve.
(a) sRABE[8] decryption
(b) sORABE decryption
(c) ciphertext size
(d) transform ciphertext size
Figure 4: (a) sRABE[8] decryption time, (b) sORABE decryption time, (c)ciphertext size and (d)transform ciphertext size for different numbers of attributes.
6.2.2 Fraud Proof Generation and Verification
We conducted the experiment and performance analysis on fraud proof generation and verification. The experiment setup and machine are the same as Section 6.2.1. The proving time and verifying time of fraud proof are about 0.87ms and 1.49ms in SS512 curve and 11.81ms and 21.40ms in SS1024 curve respectively as shown in Fig 5. It is important to note that fraud proofs are only required to be generated and verified during the dispute phase. Therefore, it does not constitute bottleneck for the system under our optimistic assumption.
(a) prove
(b) verify
Figure 5: (a) prove time and (b) verify time of fraud proof for different numbers of attributes.
(a) publishState
(b) publishTag
(c) createTask
(d) submitResult
(e) publishFraudProof
(f) publishVerifyResult
Figure 6: Gas usage of (a) publishState, (b) publishTag, (c) createTask, (d) submitResult, (e) publishFraudProof and (f) publishVerifyResult for different numbers of attributes.
6.2.3 Gas Usage
We implemented our BA-ORABE smart contract described in Section 4.4 in solidity and ran the experiment on Lenovo Thinkbook14 G5+ IRH with Intel Core i7-13700H CPU and 32GB RAM. We deploy the BA-ORABE smart contract on a locally set up Ethereum test network powered by Hardhat [^1] https://hardhat.org/, and invoke each contract function 100 times to get the average gas usage. The result is shown in Fig 6. All the functions maintain constant gas usage as the number of attributes increases. Meanwhile, except for SubmitResult and PublishFraudProof, the other functions maintain the same gas usage regardless of which curve is used. SubmitResult and PublishFraudProof cost around 2$\times$ gas in curve SS1024 than in SS512, due to the increase of the size of group $\mathbb{G}$ and $\mathbb{G}_{T}$ elements. Note that PublishFraudProof is seldom invoked under our optimistic assumption.
7 Conclusions
In this paper, we proposed BA-ORABE, a fully auditable registered attribute-based encryption scheme with reliable outsourced decryption based on blockchain, and then presented a formal security proof of its static security, verifiability, and exemptibility. Our proposed scheme achieves verifiability through a verifiable tag mechanism and exemptibility through NIZK proof. Furthermore, We extended fully auditable property and realized fairness and decentralized outsourcing based on public blockchain. Finally, we compared the theoretical complexity of BA-ORABE with current existing schemes and implemented and evaluated it on Ethereum to demonstrate efficiency and feasibility.
Funding
This work was supported by National Natural Science Foundation of China [Grant No. 62272107]; and the Innovation Action Plan of Shanghai Science and Technology [Grant No. 21511102200]; and the Key R & D Program of Guangdong Province [Grant No. 2020B0101090001].
[…]
The authors would like to thank Shuangjun Zhang, Yang Xu, and Jiamian Yan for helpful discussions on this work.
Data Availability Statement
No new data were generated or analyzed in support of this research.

### v2

@@
 A general approach to address this issue is to outsource decryption to a decryption cloud service (DCS).
-In this work, we propose BA-ORABE, the first fully auditable registered ABE with reliable outsource decryption scheme based on blockchain.
+In this work, we propose BA-ORABE, the first fully auditable registered ABE with reliable outsourced decryption scheme based on blockchain.
 First, we utilize a verifiable tag mechanism to achieve verifiability of ciphertext transformation, and the exemptibility which enables the honest DCS to escape from wrong claims is guaranteed by zero knowledge fraud proof under optimistic assumption.
 Additionally, our system achieves fairness and decentralized outsourcing to protect the interests of all parties and the registration and outsourcing process are transparent and fully auditable through blockchain.
-Finally, we give formal security analysis and implement and evaluate our scheme on Ethereum to demonstrate its feasibility and efficiency.
+Finally, we give security analysis, implement and evaluate our scheme on Ethereum to demonstrate its feasibility and efficiency, and show its advantages in real application of decentralized finance.
 keywords:
@@
 Nowadays, both individuals and enterprises are increasingly reliant on cloud storage platforms for uploading and disseminating data.
-However, given that cloud service providers (CSPs) may not always be deemed fully reliable and considering the high sensitivity of much of the data stored in the cloud, ensuring access control within cloud storage is paramount.
+However, since cloud service providers(CSPs) may not always be perceived as fully reliable, and given the high sensitivity of much of the data stored in the cloud, ensuring access control in cloud storage is crucial.
 In public cloud storage, ciphertext-policy attribute-based encryption(CP-ABE) provides an effective approach to achieve access control[1].
@@
 Although a malicious key curator could register attributes that allow the decryption of all ciphertexts, such behavior can always be detected by an external auditor.
-By utilizing a decentralized, immutable ledger, blockchain[5] can provide auditable records of all operations performed by the key curator.
+By utilizing a decentralized and immutable ledger, blockchain[5] can provide auditable records of all operations performed by the key curator.
 This ensures that any malicious behavior, such as the registration of unauthorized attributes or wrong computation, can be detected by public verifier.
@@
 Many verifiable OABE schemes have been proposed to address this issue on ABE and several OABE schemes[16, 17] achieve exemptibility but outsource heavy computation to blockchain, leading to high gas fees and performance issues.
-To the best of our knowledge, there is no outsource decryption scheme currently existing for registered ABE.
+To the best of our knowledge, there is no outsourced decryption scheme currently existing for registered ABE.
 This gap motivates our work to develop the first outsourced decryption scheme for registered ABE, which aims to reduce decryption overhead while ensuring security, verifiability, exemptibility, and fairness based on blockchain.
@@
 Besides, our scheme also achieves fairness and decentralized outsourcing, and the payment can be processed by cryptocurrency based on blockchain.
-  • We implement a concrete auditable registered ABE scheme with reliable and payable outsourced decryption and evaluate feasibility and performance on Ethereum.
+  • We implement a concrete auditable registered ABE scheme with reliable and payable outsourced decryption, evaluate its feasibility and performance on Ethereum, and show its application in decentralized finance(DeFi).
 1.2 Related Work
-1.2.1 Outsource Decryption Attribute Based Encryption
+1.2.1 Outsourced Decryption of Attribute-Based Encryption
 Attribute-Based Encryption (ABE) was first introduced as fuzzy identity-based encryption by Sahai and Waters in [18].
@@
 [19] extended it to accommodate any monotonic access policy and established a more stronger security framework compared to [11].
+But the exemptibility they claimed is not achieved since the user may reveal an incorrect private key and accuse the decryption cloud server falsely.
 None of the aforementioned schemes can achieve both exemptibility and fairness.
@@
 Furthermore, due to the lack of support for pairing computation in native Ethereum[20], this scheme is incompatible with Ethereum platform.
-1.2.2 Registered Attribute Based Encryption
+1.2.2 Registered Attribute-Based Encryption
 Registered Attribute-Based Encryption has recently been introduced as a significant extension of ABE, aimed at addressing the key escrow problem [4].
@@
 Therefore, we build upon their construction to develop an outsourced decryption scheme aimed at minimizing the decryption burden for users while achieving verifiability, exemptibility, and fairness.
+1.2.3 Applications of Optimistic Fraud Proof
+The fraud proof operates under an optimistic assumption, defaulting to accept statement unless proven otherwise.
+It includes a challenge window, during which challengers may initiate disputes to demonstrate the statement’s invalidity.
+In contrast, the validity proof offers evidence confirming the correctness of the statement, with acceptance granted only when its accuracy is definitively established under a pessimistic assumption.
+Optimistic rollups[27],[28],[29],[30],[31],[32], a key solution for Ethereum’s scalability, assumes that most transactions are valid.
+This assumption facilitates faster and more efficient transaction processing.
+To guard against invalid transactions, fraud proofs are utilized to identify and dispute inconsistencies.
+Many projects[27],[28] implement multi-round fraud proofs using the interactive bisection protocol[33], while others[31],[32] adopt zero-knowledge proofs to enable single-round fraud proofs, thereby reducing the challenge window.
+Beyond optimistic rollups, this fraud-proof mechanism is widely applied in other blockchain contexts, such as the Lightning Network[34], Plasma[35], and distributed key generation protocols[36].
 1.3 Organization
@@
 In Section 6, we implement the proposed concrete scheme on Ethereum to evaluate its performance.
+In Section 7, we describe the application of our scheme in DeFi.
 Finally, this paper is concluded in Section 8.
@@
 If $\mathbf{pk}_{u}$ is valid, KC computes and updates auxiliary data on IPFS, and publishes new state hash $\mathbf{hash(aux)}$ and public key hash $\mathbf{hash(pk_{u})}$ on blockchain.
-  • DO encrypts sensitive data using RABE under Key Encapsulation Mechanism(KEM) and sends ciphertext to IPFS.
+  • DO encrypts sensitive data using registered ABE under Key Encapsulation Mechanism(KEM) and sends ciphertext to IPFS.
   • DO publishes the associated ciphertext tag to blockchain so that DU can check the correctness of transform ciphertext with the tag.
@@
   • Smart contract will return DU’s deposit if the fraud proof is valid.
-Note that the dispute phase and public verifier are needed only when one player crashes or attempts to cheat under our optimistic assumption.
+Note that the optimistic assumption supposes that both parties are honest unless proven otherwise.
+The dispute phase and public verifier are needed only when one player crashes or attempts to cheat under our optimistic assumption.
 3.2 Threat Model
@@
 Considering that DUs intend to propose invalid fraud proof in order not to pay, we assume DU to be malicious as well.
-3.3 Security Definition
+3.3 Security Requirement
   • Static Security.
-Let $\prod_{sORABE}=(\textbf{Setup},\textbf{KeyGen},\textbf{IsValid},\textbf{Aggregate},\textbf{Encrypt},$ Transform, $\textbf{Decrypt}_{out})$ be a slotted outsource decryption registered ABE scheme with attribute universe $\mathcal{U}$ and policy space $\mathcal{P}$.
+Let $\prod_{sORABE}=(\textbf{Setup},\textbf{KeyGen},\textbf{IsValid},\textbf{Aggregate},\textbf{Encrypt},$ Transform, $\textbf{Decrypt}_{out})$ be a slotted outsourced decryption registered ABE scheme with attribute universe $\mathcal{U}$ and policy space $\mathcal{P}$.
 For a security parameter $\lambda$, an adversary $\mathcal{A}$, and a bit $b\in\{0,1\}$, we define the following game $Game_{sec}$ between a challenger C and an adversary $\mathcal{A}$:
@@
     Definition 3.1.
-    A Registered ABE scheme with outsourced decryption is secure if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
+    A registered ABE scheme with outsourced decryption is secure if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game:
     $\textbf{Pr}[b^{\prime}=b]-\frac{1}{2}\leq negl(\lambda)$
-  • Verifiability of outsourced decryption.
-Verifiability guarantees that the correctness of transform ciphertext can be efficiently checked.
-It is described by a game $Game_{ver}$ between a challenger C and an adversary $\mathcal{A}$.
-$Game_{ver}$ consists of the following steps:
-    • Setup: C runs Setup gives crs to $\mathcal{A}$.
-    • Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to C, C runs RegPK and gives mpk to $\mathcal{A}$.
-    • Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},sk^{*},CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}\}$, where S* satisfies $(M^{*},\rho^{*})$, $CT^{*}$ is an original ciphertext, $sk^{*}$ is the secret key of $\mathcal{A}$ and $CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}$ are two transformed ciphertexts of $CT^{*}$.
-$\mathcal{A}$ wins the game if:
-      $Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq\perp\cap$
-      $Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})\neq\perp\cap$
-      $Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})$
-    The advantage of the adversary in $Game_{ver}$ is defined as $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{ver}]$.
-    Definition 3.2.
-    A Registered ABE scheme with outsourced decryption is verifiable if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
-    $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{ver}]\leq negl(\lambda)$
-  • Exemptibility of outsourced decryption.
-Exemptibility guarantees that the user can not falsely accuse DCS of returning wrong transform ciphertext if it is not in the case.
-It is described by a game $Game_{exe}$ between a challenger C and an adversary $\mathcal{A}$.
-$Game_{exe}$ consists of the following steps:
-    • Setup: C runs Setup and gives crs to $\mathcal{A}$.
-    • Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to C, C runs RegPK and gives mpk and aux to $\mathcal{A}$.
-    • Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},CT^{*^{\prime}},hsk*,(sk_{1}^{*},\pi_{1}),(sk_{2}^{*},\pi_{2}),T_{i}^{*}\}$, where S* satisfies $(M^{*},\rho^{*})$, $CT^{*}$ is an original ciphertext, $CT^{*^{\prime}}$ is the transform ciphertext of $CT^{*}$, $hsk*,sk_{1}^{*},sk_{2}^{*}$ are the helper decryption key and secret keys, $T_{i}^{*}=g^{sk^{*}}$ is part of public key of $\mathcal{A}$ and $\pi_{1},\pi_{2}$ are the corresponding NIZK proof for secret keys.
-$\mathcal{A}$ wins the game if:
-      $hsk^{*}\leftarrow Update(crs,aux,pk^{*})\cap$
-      $CT^{*^{\prime}}\leftarrow Transform(hsk^{*},CT^{*})\cap$
-      $1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{1}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*}))\cap$
-      $1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{2}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*}))\cap$
-      $Decrypt_{out}(sk_{1}^{*},CT^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk_{2}^{*},CT^{*^{\prime}},CT^{*})$
-    The advantage of the adversary in $Game_{exe}$ is defined as $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{exe}]$.
-    Definition 3.3.
-    A Registered ABE scheme with outsourced decryption is exempt if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
-    $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{exe}]\leq negl(\lambda)$
+  • Verifiability of outsourced decryption. The verifiability guarantees that the correctness of transform ciphertext can be efficiently checked.
+  • Exemptibility of outsourced decryption. The exemptibility guarantees that the user can not falsely accuse DCS of returning wrong transform ciphertext through fraud proof if it is not in the case.
+  • Fairness of outsourced decryption. The fairness property in our work guarantees that:
+    • DCS gets paid if and only if it has returned a correct result.
+    • DU must pay for the correct result and can challenge wrong result through fraud proof.
+    Fairness incentivizes each party to perform the protocol honestly.
 4 Construction
@@
 Then we describe a NIZK fraud proof mechanism under optimistic assumption to achieve exemptibility.
-Finally, we present the concrete smart contract design for our payable outsourced decryption scheme.
+Finally, we present the concrete smart contract design for our auditable and fair outsourced decryption scheme.
 4.1 Verifiable Outsourced Decryption Slotted Registered ABE
@@
     $\hat{U}_{w}=\prod_{j\in[L]:w\not\in S_{j}}U_{j}\quad\text{and}\quad\hat{W}_{i,w}=\prod_{j\neq i:w\not\in S_{j}}W_{f(i,j)}.$
-    Finally, it selects three hash functions $H_{0},H_{1},H_{2}$ and outputs the master public key mpk and the slot-specific helper decryption keys $hsk_{i}$ where
+    Finally, it selects three collision-resistant hash functions $H_{0},H_{1},H_{2}$ and outputs the master public key mpk and the slot-specific helper decryption keys $hsk_{i}$ where
     $mpk=(\mathcal{G},Z,g,h,\hat{T},\hat{U}_{w},H_{0},H_{1},H_{2})$
@@
     $ct=((\textbf{M},\rho),C,C_{1},C_{2},\{C_{3,k},C_{4,k}\}_{k\in[K]},C_{5},tag)$
-  • $\textbf{Transform}(hsk,ct)$rightarrow ct’: On input the ciphertext $ct$ and the helper decryption key $hsk$. The transform algorithm proceeds as follows:
+  • $\textbf{Transform}(hsk,ct)$rightarrow ct’: On input the ciphertext $ct$ and the helper decryption key $hsk$.
 If the set of attributes $S_{i}$ is not authorized by $(\textbf{M},\rho)$, then the transform algorithm outputs $\perp$.
@@
 In this work, we make a trade-off and prioritize the security of native Ethereum.
-Note that in the registered ABE scheme, this role can also be played by KC, who is semi-honest and transparent.
+Note that in the registered ABE scheme, this public verifier role can also be played by KC, who is semi-honest and transparent.
+Figure 3: Contract design for BA-ORABE.
 4.4 Smart Contract
@@
 Transfer($U$, $x$) is a function transferring cryptocurrency of amount $x$ to user $U$.
-Figure 3: Contract design for BA-ORABE.
 5 Security Analysis
@@
 Theorem 5.1.
-Suppose the construction 4.3 of [8] is statically secure, then our sRABE with outsource decryption scheme 4.1 is also statically secure under KEM setting.
+Suppose the construction 4.3 of [8] is statically secure, then our sRABE with outsourced decryption scheme 4.1 is also statically secure under KEM setting.
 Proof 5.2.
@@
 The simulator $\mathcal{B}$ got crs from the challenger $\mathcal{C}$ and returns it to the adversary $\mathcal{A}$.
-Query. The simulator $\mathcal{B}$ answers the adversary $\mathcal{A}$’s key-generation queries on a non-corrupted slot index $i\in N$ with a set of attributes S. follows:
+Query. The simulator $\mathcal{B}$ answers the adversary $\mathcal{A}$’s key-generation queries on a non-corrupted slot index $i\in N$ with a set of attributes S as follows:
   • The simulator $\mathcal{B}$ passes it to the challenger $\mathcal{C}$ and $\mathcal{C}$ starts by incrementing the counter $ctr=ctr+1$ and samples $pk_{ctr},sk_{ctr}\leftarrow KeyGen(1^{\lambda})$.
@@
 Challenge.
-$\mathcal{A}$ submits a encapsulated key pair $(K_{0}^{*},K_{1}^{*})\in\{0,1\}^{2\times k}$ to $\mathcal{B}$, and $\mathcal{B}$ sends them to the challenger $\mathcal{C}$.
+$\mathcal{A}$ submits a plaintext pair $(K_{0}^{*},K_{1}^{*})\in\{0,1\}^{2\times k}$ to $\mathcal{B}$, and $\mathcal{B}$ sends them to the challenger $\mathcal{C}$.
 The challenger $\mathcal{C}$ generate the challenge ciphertext pair $(CT_{0}^{*},CT_{1}^{*})$ corresponds to $(K_{0}^{*},K_{1}^{*})$ and sends it to $\mathcal{B}$.
@@
 Theorem 5.3.
-Suppose our sRABE with outsource decryption scheme 4.1 is statically secure, then our RABE with outsource decryption scheme 4.2 is also statically secure.
+Suppose our slotted registered ABE with outsourced decryption scheme 4.1 is statically secure, then our registered ABE with outsourced decryption scheme 4.2 is also statically secure.
 Proof 5.4.
@@
 Theorem 5.5.
-Assume that $H_{0}$, $H_{2}$ are collision-resistant hash functions, the RABE scheme with outsourced decryption is verifiable.
+Assume that $H_{0}^{*}$, $H_{2}^{*}$ are collision-resistant hash functions, the registered ABE scheme with outsourced decryption is verifiable.
 Proof 5.6.
-Assume that there exists an adversary $\mathcal{A}$ that can break the verifiability of outsourced decryption, then we construct a simulator $\mathcal{S}$ can break the collision-resistance of hash functions $H_{0},H_{2}$.
-Given two challenging hash functions $H_{0}^{*},H_{2}^{*}$.
-$\mathcal{S}$ interacts with $\mathcal{A}$ as follows:
-Setup: $\mathcal{S}$ runs Setup gives crs to $\mathcal{A}$.
-Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to $\mathcal{S}$, $\mathcal{S}$ runs RegPK and gives mpk to $\mathcal{A}$.
-Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},sk^{*},CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}\}$, where S* satisfies $(M^{*},\rho^{*})$.
-$\mathcal{A}$ wins the game if:
-$Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq\perp\cap$
-$Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})\neq\perp\cap$
-$Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})$
-And then $\mathcal{S}$ can get:
+The verifiability is broken if DCS can return two different transform ciphertexts to match original tag. Then $\mathcal{S}$ can get:
 $H_{2}^{*}(H_{0}^{*}(\mu_{1}^{*})||C)=tag=H_{2}^{*}(H_{0}^{*}(\mu_{2}^{*})||C),\mu_{1}^{*}\neq\mu_{2}^{*}.$
 If $H_{0}^{*}(\mu_{1}^{*})=H_{0}^{*}(\mu_{2}^{*})$, then $\mathcal{S}$ can break the collision-resistance of $H_{0}^{*}$, else $\mathcal{S}$ can break the collision-resistance of $H_{2}^{*}$.
-Table 1: Comparison of Pairing-Based RABE Schemes
+5.3 Exemptibility
+Theorem 5.7.
+Suppose that the soundness of $\prod_{DLEQ-NIZK}$ protocol holds, the registered ABE scheme with outsourced decryption is exempt.
+Proof 5.8.
+The exemptibility is broken if DU can return two different valid verification keys $vk_{1}$ and $vk_{2}$ in fraud proof. Then $\mathcal{S}$ can get:
+$1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{1},(C_{2}^{\prime},vk_{1},g,T_{i}))$
+$1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{2},(C_{2}^{\prime},vk_{2},g,T_{i}))$
+where
+$vk_{1}\neq vk_{2}.$
+Thus $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}$ protocol.
+5.4 Fairness
+Theorem 5.9.
+The registered ABE scheme with outsourced decryption based on Ethereum keeps fairness.
+Proof 5.10.
+The fairness of our scheme ensures that the DCS gets paid if and only if it returns a correct result.
+The DU must deposit tokens into the smart contract to submit the outsourced decryption task.
+If the DCS submits the correct result, it will receive the reward at the end of the challenge window, which is ensured by exemptibility.
+Otherwise, the malicious behavior will be detected by DU, which is ensured by verifiability, and DCS get nothing but pay for the transaction gas fee for malicious behavior.
+Thus the proposed scheme achieves the fairness property as long as the verifiability, the exemptibility, and Ethereum consensus protocol is guaranteed.
+Table 1: Comparison of Pairing-Based Registered ABE Schemes
 […]
@@
 A scheme is statically secure if the adversary is not allowed to make any corruption queries in the security game Definition 3.1 and that it is adaptively secure otherwise.
-5.3 Exemptibility
-Theorem 5.7.
-Suppose that the soundness of $\prod_{DLEQ-NIZK}$ protocol holds, the RABE scheme with outsourced decryption is exempt.
-Proof 5.8.
-Assume that there exists an adversary $\mathcal{A}$ that can break the exemptibility of outsourced decryption, then we construct a simulator $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}$ protocol.
-Given challenging protocol $\prod_{DLEQ-NIZK}^{*}$.
-$\mathcal{S}$ interacts with $\mathcal{A}$ as follows:
-Setup: $\mathcal{S}$ runs Setup gives crs to $\mathcal{A}$.
-Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to $\mathcal{S}$, $\mathcal{S}$ runs RegPK and gives mpk and aux to $\mathcal{A}$.
-Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},CT^{*^{\prime}},hsk*,(sk_{1}^{*},\pi_{1}),(sk_{2}^{*},\pi_{2}),T_{i}^{*}\}$, where S* satisfies $(M^{*},\rho^{*})$.
-$\mathcal{A}$ wins the game if:
-$hsk^{*}\leftarrow Update(crs,aux,pk^{*})\cap$
-$CT^{*^{\prime}}\leftarrow Transform(hsk^{*},CT^{*})\cap$
-$1\leftarrow\prod_{DLEQ-NIZK}^{*}.\textbf{Verify}(\pi_{1}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*}))\cap$ $1\leftarrow\prod_{DLEQ-NIZK}^{*}.\textbf{Verify}(\pi_{2}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*}))\cap$ $Decrypt_{out}(sk_{1}^{*},CT^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk_{2}^{*},CT^{*^{\prime}},CT^{*}).$ And then $\mathcal{S}$ can get
-$sk_{1}^{*}\neq sk_{2}^{*}\ \cap$
-$(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*})\in R_{DLEQ}\ \cap$
-$(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*})\in R_{DLEQ}$
-Thus $\mathcal{S}$ can break the soundness of $\prod_{DLEQ-NIZK}^{*}$ protocol.
-5.4 Fairness
-The fairness of our scheme ensures that the DCS gets paid if and only if it returns a correct result.
-The DU must deposit tokens into the smart contract to submit the outsourced decryption task.
-If the DCS submits the correct result, it will receive the reward at the end of the challenge window, which is ensured by exemptibility.
-Otherwise, DU will send valid fraud proof and get the deposit back, and DCS get nothing but pay for the transaction gas fee for malicious behavior.
-Thus the proposed scheme achieves the fairness property as long as the exemptibility and Ethereum consensus protocol is guaranteed.
-Table 2: Comparison With Outsourced Decryption ABE Schemes
-Schemes confidentiality verifiability exemptibility fairness decentralized outsourcing
-[11]
-[12]
-[13]
-[14]
-[19]
-[17]
-Our Scheme
 6 Performance and Evaluation
@@
 And the parameter sizes are better than other schemes.
-Besides, our outsourced decryption significantly reduces the final decryption cost to a constant level, requiring only one group exponentiation operation and one group multiplication operation.
-Table 2 shows the functional comparison of the previous OABE schemes and our scheme in terms of confidentiality, verifiability, exemptibility, fairness, and decentralized outsourcing.
-Almost all schemes achieve confidentiality and verifiability.
-Only our scheme and [17] achieves exemptibility, fairness, and decentralized outsourcing property.
+Besides, our outsourced decryption significantly reduces the final decryption cost to a constant level without any pairing operation, requiring only one group exponentiation operation.
+Table 2 shows the comparison of the previous OABE schemes and our scheme in terms of authority, verifiability, exemptibility, fairness, decentralized outsourcing, and final decryption cost.
+Our scheme is the only one that does not rely on a trusted authority and achieves the lowest decryption overhead among all the compared schemes.
+While most schemes ensure verifiability, only our scheme and [17] provide exemptibility, fairness, and decentralized outsourcing capabilities.
 6.2 Experimental Evaluation
@@
 As we use the slotted registered ABE to registered ABE transformation[4], which constructs registered ABE through multiple slotted registered ABE instances, We only compared the basic slotted scheme here.
-We conducted experiment and performance analysis on the computation cost comparison between sRABE scheme in [8] and our sRABE with outsourced decryption scheme.
+We conducted experiment and performance analysis on the computation cost comparison between sRABE scheme in [8] and our sRABE with outsourced decryption scheme, excluding the computational cost of symmetric encryption for consistency.
 We implemented them in Python language within the Charm framework.
@@
 For each algorithm, we repeat 1,000 times to get the average execution time.
-As shown in Fig 4(a), the decryption time of sRABE scheme in [8] increases linearly with the numbers of attributes while our scheme with outsource decryption has constant decryption time and is 108$\times$-1585$\times$ in SS512 curve and 175$\times$-2647$\times$ in SS1024 curve lower than [8] in attribute numbers from 5 to 100, which is shown in Fig 4(b).
-The original and transform ciphertext size of our scheme are shown in Fig 4(c) and Fig 4(d) respectively.
-Note that we only consider RABE ciphertext and the symmetric encryption ciphertext is not included.
-The original ciphertext linearly increases with the numbers of attributes like [8] and the transform ciphertext contains only two $\mathbb{G}_{T}$ elements, which is about 0.25kb in SS512 curve and 0.5kb in SS1024 curve.
 (a) sRABE[8] decryption
@@
 Figure 4: (a) sRABE[8] decryption time, (b) sORABE decryption time, (c) ciphertext size and (d) transform ciphertext size for different numbers of attributes.
+As shown in Fig 4(a), the decryption time of sRABE scheme in [8] increases linearly with the numbers of attributes while our scheme with outsourced decryption has constant decryption time and is 108$\times$-1585$\times$ in SS512 curve and 175$\times$-2647$\times$ in SS1024 curve lower than [8] in attribute numbers from 5 to 100, which is shown in Fig 4(b).
+The original and transform ciphertext size of our scheme are shown in Fig 4(c) and Fig 4(d) respectively.
+Note that we only consider registered ABE ciphertext and the symmetric encryption ciphertext is not included.
+The original ciphertext linearly increases with the numbers of attributes like [8] and the transform ciphertext contains only two $\mathbb{G}_{T}$ elements, which is about 0.25kb in SS512 curve and 0.5kb in SS1024 curve.
+Table 2: Comparison With Outsourced Decryption ABE Schemes
+Schemes Authority-Free Verifiability Exemptibility Fairness Decentralized Outsourcing Decryption Cost
+[…]
+[12] $\mathbb{E}_{T}$
+[13] $\mathbb{E}_{T}$
+[…]
+[…]
+[17] $\mathbb{E}_{T}$
+Our Scheme $\mathbb{E}_{T}$
 6.2.2 Fraud Proof Generation and Verification
-We conducted the experiment and performance analysis on fraud proof generation and verification.
+We conducted the experiment on fraud proof generation and verification.
 The experiment setup and machine are the same as Section 6.2.1.
@@
 Figure 5: (a) prove time and (b) verify time of fraud proof for different numbers of attributes.
-(a) publishState
-(b) publishTag
-(c) createTask
-(d) submitResult
-(e) publishFraudProof
-(f) publishVerifyResult
-Figure 6: Gas usage of (a) publishState, (b) publishTag, (c) createTask, (d) submitResult, (e) publishFraudProof and (f) publishVerifyResult for different numbers of attributes.
 6.2.3 Gas Usage
@@
 Note that PublishFraudProof is seldom invoked under our optimistic assumption.
-7 Conclusions
-In this paper, we proposed BA-ORABE, a fully auditable registered attribute-based encryption scheme with reliable outsourced decryption based on blockchain, and then presented a formal security proof of its static security, verifiability, and exemptibility.
-Our proposed scheme achieves verifiability through a verifiable tag mechanism and exemptibility through NIZK proof.
+Figure 6: Gas usage of different contract functions
+7 Application
+Decentralized Finance(DeFi) [39] signifies a fundamental shift from traditional banking and financial services, providing a blockchain-based alternative that eliminates the reliance on central authorities.
+By utilizing smart contracts on the Ethereum blockchain, DeFi allows users to lend, borrow, trade assets, and earn interest within a permissionless and transparent ecosystem.
+Applying ABE in DeFi can provide users with fine-grained access control for transactions, enhancing the security of digital asset trading in decentralized exchanges.
+Ma et al.[40] proposed a scheme based on traditional ABE where users define attributes such as asset type, amount, exchange rate, and time, and set diverse access control policies.
+Only transaction counterparts that satisfy the user’s specified policies are allowed to engage in the trade, thereby enhancing transaction security and ensuring compliance with user preferences in decentralized exchanges.
+However, traditional ABE schemes rely on one or multiple trusted authorities, which still poses certain security concern when applied to DeFi.
+In contrast, registered ABE eliminates the need for any trusted authority, and its transparent process makes it more suitable for application in DeFi.
+Our auditable scheme further enhances the security of the system.
+Moreover, due to the widespread use of mobile devices for transactions, the high decryption overhead in ABE presents a significant challenge.
+The outsourced decryption registered ABE scheme effectively addresses this issue.
+In summary, the auditable registered ABE scheme with outsourced decryption proposed in this work provides significant advantages for transaction control in DeFi.
+8 Conclusion
+In this paper, we proposed BA-ORABE, a fully auditable registered attribute-based encryption scheme with reliable outsourced decryption based on blockchain, and then presented security proof of its static security, verifiability, and exemptibility.
+Our proposed scheme achieves verifiability through verifiable tag mechanism and exemptibility through NIZK proof.
 Furthermore, We extended fully auditable property and realized fairness and decentralized outsourcing based on public blockchain.
 Finally, we compared the theoretical complexity of BA-ORABE with current existing schemes and implemented and evaluated it on Ethereum to demonstrate efficiency and feasibility.
+We also showed that our scheme provides significant advantages in DeFi application, which could be explored further in future work.
 Funding